# Sensing and Steering Stereotypes: Extracting and Applying Gender Representation Vectors in LLMs

### v2

@@ Abstract
Large language models (LLMs) are known to perpetuate stereotypes and exhibit biases. Various strategies have been proposed to mitigate these biases, but most work studies biases as a black-box problem without considering how concepts are represented within the model. We adapt techniques from representation engineering to study how the concept of “gender” is represented within LLMs. We introduce a new method that extracts concept representations via probability weighting without labeled data and efficiently selects a steering vector for measuring and manipulating the model’s representation. We develop a projection-based method that enables precise steering of model predictions and demonstrate its effectiveness in mitigating gender bias in LLMs. [^1] Our code is available at: https://github.com/hannahxchen/gender-bias-steering
Sensing and Steering Stereotypes:
Extracting and Applying Gender Representation Vectors in LLMs
@@ 1 Introduction
While much research has explored gender bias in LLMs through a black-box approach, less attention has been paid to how these biases arise from the model’s internal workings. Recent work on representation engineering provides insights into varied abstract features within the internal representations of LLMs (Zou et al., 2023), such as sentiment (Tigges et al., 2023), spatiotemporal information (Gurnee and Tegmark, 2024), and true/false statements (Marks and Tegmark, 2024). Several studies have also demonstrated promising results in effectively controlling model behaviors by modifying their feature representations (Turner et al., 2023; Rimsky et al., 2024; Arditi et al., 2024).
In this work, we leverage activation steering (also known as activation engineering), to study how the concept of gender is encoded in the internal representations of LLMs affects their predictions and how we can manipulate internal representations to mitigate biases at inference time.
@@
Contributions. We propose a novel method that extracts linear representations from LLMs for steering model predictions associated with a given concept (Section 3). While existing methods for computing steering vectors rely on labeled data, we compute them using probability weighting without explicit data annotations. In addition, we introduce metrics to efficiently select a steering vector without exhaustive searches as was required by most previous methods. We show that steering vectors produced by our method exhibit a higher correlation with gender bias in model outputs than the prevailing difference-in-means method (Section 3.4). We then present an approach for applying steering vectors to provide precise control over the internal representation (Section 4). We demonstrate the effectiveness of our steering vectors and method for applying them in reducing gender bias on the in-distribution task (Section 4.2) and its potential to generalize to other application tasks (Section 4.3). Finally, we explore the generalization of our method for controlling bias associated with other protected attributes (Section 5).
@@ 3.1 Extracting Candidate Vectors
Let $A$ and $B$ denote two contrasting concepts (e.g., femaleness and maleness) each of which can be identified by an associated set of tokens. We measure the extent of $A$ and $B$ presented in a model for an input prompt $x\in\mathcal{D}$ based on its prediction output. We define the disparity score between the two concepts for an input $x$ as:
@@
where $P_{x}(A)$ is the probability of predicting concept $A$ in the last token position output of $x$, aggregated over tokens for $A$. The disparity score indicates how likely an input would trigger the model to predict one concept over another in the next token prediction.
@@
In contrast to difference-in-means, which computes the activation mean difference between $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, we incorporate neutral prompts with probability weighting to filter out signals unrelated to the target concepts. This allows the vector to capture a better representation of $A$ and $B$.
@@ 3.3 Experimental Setup
Our prompts ask the model to respond with the gender indicated in the given sentence, followed by a sentence from the dataset. Since some models do not directly respond with a gender-related token, we add an output prefix to guide the model to produce more relevant outputs in the next token prediction. For each gender concept, we randomly sample 800 prompts that satisfy the requirements of Section 3.1 for extracting the candidate vectors. The number of neutral prompts varies by model, but we subsample them if the size is larger than either the set of gendered prompts. We set the default score threshold $\delta$ to 0.05, but compare results using different $\delta$ values in Section E.2. Appendix A provides more details, including the gender tokens used for computing the disparity scores.
@@
Figure 3: Disparity scores $s_{x}$ and scalar projections of each input from the validation set. The first and third columns show the baseline measured before intervention. The second and fourth columns illustrate the change in disparity scores by overlaying the results after debiasing from the left figures. The projections (x-axis) of all datapoints are measured before intervention. We perform intervention at the layer where the vector has the lowest RMSE.
@@ 3.4 Results
Figure 3 (first and third columns) compares the disparity scores and scalar projections measured for each input prompt with the steering vector selected at the optimal layer. Ideally, the projections should align closely with the green dashed line in the figure, reflecting a positive correlation with the disparity scores measured in model outputs. Our proposed method WMD yields a better correlation with the disparity scores, where inputs with a higher disparity show a larger projection value, as measured by the selected steering vector. It also reflects the degree of disparities more equally in both female and male directions. While MD does capture the gender representations to some extent, it poorly reflects with inputs more associated with the maleness concept where $s_{x}<0$, as shown in 3(b) for Qwen-7B model. For some of these inputs, the projections on the steering vector indicate a higher degree of female signal. This imbalance in generalization may impact their steering performance, which we demonstrate in the next section.
Baseline MD WMD Modal Interval
@@
Llama-2-13B 0.49 29 0.81 0.28 37 0.85 0.16 [--0.33, 0.18]
Llama-3.1-8B 0.65 26 0.84 0.60 25 0.98 0.32 [--0.23, 0.15]
Ministral-8B 0.50 30 0.95 0.05 27 0.95 0.07 [--0.10, 0.12]
Mistral-Nemo-12B 0.65 35 0.89 0.08 37 0.98 0.02 [--0.32, 0.00]
Qwen-1.8B 0.53 19 0.88 0.14 19 0.88 0.14 [--0.95, 0.99]
Qwen-7B 0.51 26 0.69 0.32 29 0.88 0.12 [--0.27, 0.22]
Granite-3.1-8B 0.63 37 0.96 0.27 37 0.97 0.24 [--0.05, 0.05]
OLMo-2-7B 0.63 29 0.88 0.47 27 0.90 0.37 [--0.44, 0.16]
@@ 4.1 Intervention Method
Since a model can exhibit varied degrees of bias to different inputs and at different generation steps, we cannot achieve precise control of model behaviors by simply applying activation addition with a uniform steering coefficient (Equation 1). To obtain more precise control, we perform interventions for each input $x$ as follows:
@@
where $\lambda$ is the steering coefficient and $\hat{\bm{v}}$ is $\vec{v}$ in unit vector form. When $\lambda=0$, it performs bias mitigation by subtracting the vector projection $\mathbf{proj}_{\bm{v}}x$, which captures the extent of bias in input $x$ reflected on the steering vector $\vec{v}$. We can steer the model outputs to either concept with a coefficient value of $\lambda\not=0$. The model becomes more biased to $A$ when $\lambda>0$ and to $B$ when $\lambda<0$. We apply this operation across all token positions of $x$ but at only the layer from which $\vec{v}$ was extracted.
This formulation is similar to directional ablation proposed by Arditi et al. (2024), which also considers vector projections. However, they show that this approach, using steering vectors computed by MD, can only be used for removing a single concept (in one direction) and requires interventions across all model layers. Our proposed intervention provides a unified formulation for concept removal and steering model behaviors in either direction.
@@ 4.2 Steering for Bias Mitigation
We evaluate the effectiveness of steering vectors selected using the method described in Section 3.4 to mitigate gender bias. We apply the steering vectors with our proposed projection-based intervention with $\lambda=0$ and measure the bias score on the validation set, computed as the root mean square (RMS) of disparity score $s_{x}$.
Table 1 reports the bias scores before and after debiasing for each model. After applying the intervention, it shows a significant reduction in the bias score for all models. The intervention is particularly effective for Ministral-8B and Mistral-Nemo-12B instruction models with bias scores reduced to nearly zero. In addition, the results suggest that the projection and bias score correlation $r$ is a good indicator of the intervention performance. Models with a higher value of $r$ show a greater decrease in the bias score after intervention.
To analyze the impact of intervention on different inputs, we compare the bias score difference and the scalar projection of each input, as shown in the second and fourth columns of Figure 3. We apply the same intervention method for both steering vectors computed by MD and WMD. The projections of all data points are measured on the baseline model with no intervention. Debiasing with WMD’s steering vectors works as intended, where more biased inputs show a larger difference in their bias scores after debiasing, while less biased inputs are less affected. However, the inputs tend to be over- or under-corrected in their bias scores when using steering vectors computed by MD. As our intervention approach depends on the projection of each input, the mitigation becomes less effective when the steering vector fails to separate the bias direction or does not reflect well with model bias.
@@ 4.3.1 Evaluation Tasks
Winogenerated (Perez et al., 2023) is a human validated version of the Winogender pronoun resolution task (Rudinger et al., 2018) that is 50 times larger than the original datset. The model is asked to fill in the missing blank with a pronoun for a given sentence (e.g., “The surgeon assured the patient that __ would do the best possible job.”). The response can be either a male, female, or gender-neutral pronoun. We report the output probability produced for each gender pronoun, normalizing over all three pronoun options.
@@ 4.3.2 Results
When $\lambda=0$, we remove bias by the degree of gender signal reflected on the steering vector (Equation 4). As shown in Figure 4, when $\lambda\simeq 0$, our method effectively reduces gender bias in the model, with neutral pronouns having the highest probability, while male and female pronouns show similar but lower probabilities on average. The effect of coefficient values on the model’s outputs also aligns with the expected gender concept. A more positive $\lambda$ increases the output probability for female pronouns, whereas a more negative $\lambda$ increases it for male pronouns. The model is less likely to predict neutral pronouns when steering with a larger magnitude of $\lambda$ in either direction.
For the occupational stereotypes task, we analyze the frequency difference in job titles predicted for feminine and masculine names before and after debiasing with steering vectors. Figure 5 displays the predicted job titles in the technology and healthcare sectors with the largest gender disparities. Prior to intervention, the model exhibits the largest discrepancies in predicting “software engineer” and “product manager” in technology and “nurse” and “doctor” in healthcare. Debiasing substantially decreases the frequency gap for most common job titles, and increases the relative prediction frequency of more neutral titles, such as “healthcare professional” for masculine names.
@@
Figure 5: Difference in job title prediction frequency when prompted with feminine names compared to masculine names. The color represents the difference before and after debiasing on Qwen-1.8B. The y-axis shows the top 10 titles with the largest prediction gap.
Figure 6: Scalar projections of examples from the occupational stereotypes task, evaluated on Qwen-1.8B at the last token position. The color indicates the gender associated with the name used in the prompt.
5 Steering Racial Bias
This section explores generalizing our method to other protected attributes. We experiment with the concept of racial majority–minority groups, where the majority is White American and the minority is Black American. [^2] As with gender, race is a complex and non-binary notion that cannot be fully captured with a single dimension. We do not intend to suggest any kind of racial binary by using these categories, just select these as representative categories to enable our experiments because of the availability of data from previous linguistic experiments. We show that our proposed mitigation can be applied similarly to reduce racial bias in models.
@@ 5.2 Results
7(a) compares the disparity scores before and after mitigating racial bias with the steering vectors we found for Llama-3.1-8B and Mistral-Nemo-12B. The steering vectors for both models show a strong correlation with the disparity scores before debiasing. In 7(b), we compare the model’s output probabilities for both racial concepts when applied with different steering coefficients $\lambda$. The probabilities (as shown by the solid lines) are measured by the normalized output probabilities of white- and black-associated tokens, averaged over 200 sampled inputs. The result after debiasing in 7(a) corresponds to $\lambda=0$ in 7(b). Both models show a similar probability between white and black, which aligns with our intended goal of debiasing. The effect of the coefficient value $\lambda$ is also consistent with the desired model behavior. A higher positive value increases the probability of predicting black-associated tokens, whereas a larger negative $\lambda$ increases the probability of predicting white-associated tokens.
Our results demonstrate how our proposed method can be generalized to manipulate other protected attributes and mitigate bias in models. Additional results are provided in Section D.3.

### v3

# Unsupervised Concept Vector Extraction for Bias Control in LLMs

@@ Abstract
Large language models (LLMs) are known to perpetuate stereotypes and exhibit biases. Various strategies have been proposed to mitigate these biases, but most work studies biases as a black-box problem without considering how concepts are represented within the model. We adapt techniques from representation engineering to study how the concept of “gender” is represented within LLMs. We introduce a new method that extracts concept representations via probability weighting without labeled data and efficiently selects a steering vector for measuring and manipulating the model’s representation. We develop a projection-based method that enables precise steering of model predictions and demonstrate its effectiveness in mitigating gender bias in LLMs and show that it also generalizes to racial bias. [^1] Our code is available at: https://github.com/hannahxchen/gender-bias-steering
Unsupervised Concept Vector Extraction for Bias Control in LLMs
@@ 1 Introduction
While much research has explored gender bias in LLMs through a black-box approach, less attention has been paid to understanding how these biases arise from the model’s internal workings. Recent work on representation engineering provides insights into varied abstract features within the internal representations of LLMs (Zou et al., 2023), such as sentiment (Tigges et al., 2023), spatiotemporal information (Gurnee and Tegmark, 2024), and true/false statements (Marks and Tegmark, 2024). Several studies have also demonstrated promising results in effectively controlling model behaviors by modifying their internal feature representations (Turner et al., 2023; Rimsky et al., 2024; Arditi et al., 2024).
In this work, we leverage activation steering (also known as activation engineering), to study how the concept of gender is encoded in the internal representations of LLMs, how it affects their predictions, and how we can manipulate internal representations to mitigate biases at inference time.
@@
Contributions. We propose a novel method that extracts linear representations from LLMs for steering model predictions associated with a given concept (Section 3). Unlike prior methods that rely on labeled data to compute steering vectors, our approach uses probability weighting without explicit data annotations. In addition, we introduce metrics to efficiently select a steering vector without exhaustive searches, as was required by most previous methods. We show that steering vectors produced by our method exhibit a higher correlation with gender bias in model outputs than the prevailing difference-in-means method (Section 3.4). We then present an approach for applying steering vectors with precise control over internal representations (Section 4). We demonstrate the effectiveness of our steering vectors and our method for applying them in reducing gender bias on the in-distribution task (Section 4.2) and its potential to generalize to other application tasks (Section 4.3), without degrading models’ general capabilities (Section 4.4). Finally, we explore the generalization of our method for controlling bias associated with other protected attributes (Section 5), showing that it can be used to understand and mitigate racial biases.
@@ 3.1 Extracting Candidate Vectors
Let $A$ and $B$ denote two contrasting concepts (e.g., femaleness and maleness), each of which can be identified by an associated set of tokens. While previous methods have treated inputs as strictly binary (either $A$ or $B$), we hypothesize that they may present varying degrees of associations with each concept, as encoded in the model. We measure the association based on the model’s prediction output. For an input prompt $x\in\mathcal{D}$, we compute a model-specific disparity score, $s_{x}$, between the two concepts, defined as:
@@
where $P_{x}(A)$ is the probability of predicting concept $A$ in the last token position output of $x$, aggregated over tokens for $A$. The disparity score indicates how likely an input would be to trigger the model to predict one concept over another in the next token prediction.
@@
In contrast to difference-in-means, which only considers $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ and treats inputs with the same label identically, we incorporate neutral prompts and apply probability weighting. We assign higher weightings to inputs that encode stronger concept signals while minimizing potential noise unrelated to the target concepts. This allows us to extract vectors that capture more accurate representations of the concepts $A$ and $B$.
@@ 3.3 Experimental Setup
Our prompts ask the model to respond with the gender indicated in the given sentence, followed by a sentence from the dataset. Since some models do not directly respond with a gender-related token, we add an output prefix to guide the model to produce more relevant outputs in the next token prediction. For each gender concept, we randomly sample 800 prompts that satisfy the requirements of Section 3.1 for extracting the candidate vectors. The number of neutral prompts varies by model, but we subsample them if the size is larger than either set of gendered prompts. We set the default score threshold $\delta$ to 0.05, but compare results using different $\delta$ values in Section E.2. Appendix A provides more details, including the gender tokens used for computing the disparity scores.
@@
Figure 3: Disparity scores $s_{x}$ and scalar projections of each input from the validation set. The first and third columns show the baseline measured before intervention. The second and fourth columns illustrate the change in disparity scores by overlaying the results after steering from the left figures. The projections (x-axis) of all datapoints are measured before intervention. We perform interventions at the layer where the vector has the lowest RMSE.
@@ 3.4 Results
Figure 3 (first and third columns) compares the disparity scores and scalar projections measured for each input prompt with the steering vector selected at the optimal layer. Ideally, the projections should align closely with the green dashed line in the figure, reflecting a positive correlation with the disparity scores measured in model outputs. Our proposed method WMD yields a better correlation with the disparity scores, where inputs with a higher disparity show a larger projection value, as measured by the selected steering vector. It also reflects the degree of disparities more equally in both female and male directions. While MD captures the gender representations to some extent, it poorly reflects inputs more associated with the maleness concept where $s_{x}<0$, as shown in 3(b) for Qwen-7B model. For some of these inputs, the projections on the steering vector indicate a higher degree of female signal. This imbalance in generalization may impact the steering performance, which we demonstrate in the next section.
Baseline MD WMD
@@
Llama-2-13B 0.49 29 0.81 0.28 37 0.85 0.16
Llama-3.1-8B 0.65 26 0.84 0.60 25 0.98 0.32
Ministral-8B 0.50 30 0.95 0.05 27 0.95 0.07
Mistral-Nemo-12B 0.65 35 0.89 0.08 37 0.98 0.02
Qwen-1.8B 0.53 19 0.88 0.14 19 0.88 0.14
Qwen-7B 0.51 26 0.69 0.32 29 0.88 0.12
Granite-3.1-8B 0.63 37 0.96 0.27 37 0.97 0.24
OLMo-2-7B 0.63 29 0.88 0.47 27 0.90 0.37
@@ 4.1 Intervention Method
Since a model can display varying degrees of concept associations with different inputs and at various generation steps, we cannot achieve precise control of model behaviors by simply applying activation addition with a uniform steering coefficient (Equation 1). To obtain more precise control, we perform interventions for each input $x$ as follows:
@@
where $\lambda$ is the steering coefficient and $\hat{\bm{v}}$ is $\vec{v}$ in unit vector form. When $\lambda=0$, we subtract the activation by its vector projection $\mathbf{proj}_{\bm{v}}x$, thereby removing any signals related to either concept. To steer model outputs to one of the concepts, we apply a non-zero coefficient value. The model increases association with $A$ when $\lambda>0$ and with $B$ when $\lambda<0$. This operation is applied across all token positions of $x$ but at only the layer from which $\vec{v}$ was extracted.
Previous work has proposed interventions using vector projections. Arditi et al. (2024) apply directional ablation to remove concept representations, using steering vectors computed by MD. However, this method can only be used for removing a single concept (in one direction) and requires interventions across all model layers. Lee et al. (2025) steer model behaviors conditionally based on projections. Our proposed intervention provides a unified formulation for concept removal and steering model behaviors in either direction.
@@ 4.2 Steering for Bias Mitigation
We assess the effectiveness of steering vectors found in Section 3.4 to mitigate gender bias. Consider a steering vector $\vec{v}$, which encodes the gender concept of the model. We assume that the projection on $\vec{v}$ reflects the degree of gender signals in the model and that removing the signals can help reduce gender bias in its predictions. We perform bias mitigation using our proposed projection-based intervention (Equation 4) with $\lambda=0$. We evaluate the debiasing performance based on the bias score on the validation set, computed as the root mean square (RMS) of disparity score $s_{x}$.
Table 1 reports the bias scores before and after steering with $\lambda=0$. We apply the same intervention method for both steering vectors computed by MD and WMD. After applying the intervention, the bias score for all models shows a significant reduction. The intervention is particularly effective for Ministral-8B and Mistral-Nemo-12B instruction models with bias scores reduced to nearly zero. In addition, the results suggest that the projection and bias score correlation $r$ is a good indicator of the intervention performance. Models with a higher value of $r$ show a greater decrease in the bias score after intervention.
To analyze the impact of intervention on different inputs, we compare the change in disparity score and the scalar projection of each input, as shown in the second and fourth columns of Figure 3. The projections of all data points are measured on the baseline model with no intervention. Debiasing with WMD’s steering vectors works as intended, where more “biased” inputs show a larger difference in their disparity scores after intervention, while less “biased” inputs are less affected. However, the inputs tend to be over- or under-corrected in their disparity scores when using steering vectors computed by MD. As our intervention approach depends on the projection of each input, the mitigation becomes less effective when the steering vector fails to separate the gender direction or does not reflect well with the disparity score.
@@ 4.3.1 Evaluation Tasks
Winogenerated (Perez et al., 2023) is a human validated version of the Winogender pronoun resolution task (Rudinger et al., 2018) that is 50 times larger than the original dataset. The model is asked to fill in the missing blank with a pronoun for a given sentence (e.g., “The surgeon assured the patient that __ would do the best possible job.”). The response can be either a male, female, or gender-neutral pronoun. We report the output probability produced for each gender pronoun, normalizing over all three pronoun options.
@@ 4.3.2 Results
When $\lambda=0$, gender signals are expected to be eliminated from the model. As shown in Figure 4, the model predicts neutral pronouns with the highest probability when $\lambda\simeq 0$, while male and female pronouns have similar but lower probabilities on average. The effect of coefficient values on the model’s outputs also aligns with the expected gender concept. A more positive $\lambda$ increases the output probability for female pronouns, whereas a more negative $\lambda$ increases it for male pronouns. The model is less likely to predict neutral pronouns when steering with a larger magnitude of $\lambda$ in either direction.
For the occupational stereotypes task, we analyze the frequency difference in job titles predicted for feminine and masculine names before and after removing gender signals with steering. Figure 5 displays the predicted job titles in the technology and healthcare sectors with the largest gender disparities. Prior to intervention, the model exhibits the largest discrepancies in predicting “software engineer” and “product manager” in technology and “nurse” and “doctor” in healthcare. After intervention, the differences for these common job titles decrease substantially, and neutral titles such as “healthcare professional” are predicted more frequently for masculine names.
Figure 5: Difference in job title prediction frequency when prompted with feminine names compared to masculine names. The color represents the difference before and after removing gender signals in Qwen-1.8B when steering with $\lambda=0$. The y-axis shows the top 10 titles with the largest prediction gap.
@@
Our results suggest that the steering vectors we found capture other forms of gender associations encoded in the model, beyond gendered pronouns. Section G.1 provides several model outputs produced by steering, which show changes in gender stereotypes related to appearance, personality, hobbies, and occupations.
Figure 6: Scalar projections of inputs for the occupational stereotypes task, evaluated on Qwen-1.8B at the last token position. The color indicates the gender associated with the name used in the prompt.
Model MMLU IFEval
Acc Strict Acc Loose Acc
Qwen-7B 0.553 / 0.553 0.309 / 0.316 0.323 / 0.327
Ministral-8B 0.618 / 0.618 0.486 / 0.462 0.506 / 0.490
Llama2-13B 0.512 / 0.511 0.323 / 0.314 0.445 / 0.436
Table 2: Model performance on general language benchmarks after debiasing with gender steering vectors, followed by the baseline performance before intervention. We report MMLU by the average accuracy on the test set and IFEval by the prompt-level strict and loose accuracies.
4.4 Model Quality
We evaluate the effect of steering on the overall model capability based on two general language benchmarks: (1) MMLU, which tests the model’s knowledge and problem-solving abilities on multiple-choice questions (Hendrycks et al., 2021), and (2) IFEval, which evaluates the instruction-following ability based on a set of verifiable instructions (Zhou et al., 2023). We report the average accuracy of MMLU on the test set, which contains 14,042 questions, and the prompt-level strict and loose accuracies on IFEval, which includes 541 instructions.
Table 2 compares the performance before and after debiasing the model with the gender steering vector we found in Section 3.4. We find that debiasing with steering has little impact on the model’s overall capability. Since we debias only by vector projections with $\lambda=0$ (Equation 4), this may mean that the MMLU and IFEval inputs simply do not exhibit any gender signal ($\mathbf{proj}_{\bm{v}}x\approx 0$), as indicated by the steering vector.
Next, we evaluate the impact of steering coefficients on model performance on the two tasks. We rescaled the steering vector based on the ratio of scalar projection to disparity score measured from the validation set, so that we can simply steer between $\lambda\in[-1,1]$. (We consider the valid disparity score range to be $s_{x}\in[-1,1]$.) To reduce computational overhead, we apply steering to MMLU only for the high school subjects, which contain 3,420 questions. Figure 7 shows model performance when steering with different coefficients. Our finding suggests that increasing the coefficient in either direction does not significantly affect the overall performance on these two tasks (at least within the valid coefficient range). It is likely that the gender concept captured by our method is orthogonal to the model’s ability to solve the tasks.
Figure 7: Model performance on general language and instruction-following benchmarks when steering along the gender vector with a coefficient $\lambda\in[-1,1]$.
5 Steering Racial Concepts
So far, we have demonstrated how our method can be applied to find steering vectors for manipulating gender signals in models. This section explores generalizing our method racial majority–minority groups, where the majority is White American and the minority is Black American. [^2] As with gender, race is a complex and non-binary notion that cannot be fully captured with a single dimension. We do not intend to suggest any kind of racial binary by using these categories, just select these as representative categories to enable our experiments because of the availability of data from previous linguistic experiments. We show that our proposed mitigation can be applied similarly to reduce racial bias in models.
@@ 5.2 Results
8(a) compares the disparity scores before and after removing racial signals with the steering vectors we found for Llama-3.1-8B and Mistral-Nemo-12B. The steering vectors for both models show a strong correlation with the disparity scores before debiasing. In 8(b), we compare the model’s output probabilities for both racial concepts when applied with different steering coefficients $\lambda$. The probabilities (as shown by the solid lines) are measured by the normalized output probabilities of white- and black-associated tokens, averaged over 200 sampled inputs. The result after debiasing in 8(a) corresponds to $\lambda=0$ in 8(b). Both models show a similar probability between white and black, which aligns with our intended goal of debiasing. The effect of the coefficient value $\lambda$ is also consistent with the desired model behavior. A higher positive value increases the probability of predicting black-associated tokens, whereas a larger negative $\lambda$ increases the probability of predicting white-associated tokens.
Our results demonstrate how our proposed method can be used for controlling bias related to other protected attributes in LLMs. Additional results are provided in Section D.3.